OpenStudy (anonymous):

Given that the acceleration vector is a(t)=(-4cos(-2t))i + (-4sin(-2t))j + (4t)k, the initial velocity is v(0)=(i + k) , and the initial position vector is r(0)= i + j + k , compute v(t)=? r(t)=?

OpenStudy (anonymous):

so you know that acceleration is your second derivative of your position/vector "whatchamacall it" ... you can say r(t) correct?

OpenStudy (anonymous):

So you must integrate a(t) in which you'll have to integrate each component by itself due to it being a vector

OpenStudy (anonymous):

i have integrated already i got v(t)=(-2sin(2t)+1)i +(2cos(2t))j + (2t^2+1)k

OpenStudy (anonymous):

but j is wrong

OpenStudy (anonymous):

+C

OpenStudy (anonymous):

they said j was wrong? let me check to see what is up

OpenStudy (anonymous):

the integration of sin will give you a -cos

OpenStudy (anonymous):

so before doing any integration pull the -4 out and then make u=-2t

OpenStudy (anonymous):

so i got 2cos(-2t)

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

\[-4\int\limits_{}^{}\sin(-2t)dtj\] let u= -2t du = -2dt du/-2=dt if you pull that -2 out you get \[2\int\limits_{}^{}\sin(u)du\] \[2(-\cos(u) + C)\]

OpenStudy (anonymous):

did they get -2cos(-2t)

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

what do they have?

OpenStudy (anonymous):

its my hw question i cant figure it out

OpenStudy (anonymous):

oh well thats not the answer yet lol that's just j... you said that the integration was wrong

OpenStudy (anonymous):

after the integration you get + C correct so you need to find your constant vector

OpenStudy (anonymous):

C = to v(0)

OpenStudy (anonymous):

and it 0 according to v(0)

OpenStudy (anonymous):

so after integration of your acceleration vector did you get \[2\sin(-2t)i-2\cos(-2t)j+2t^2k\]

OpenStudy (anonymous):

yea i got that

OpenStudy (anonymous):

and only j was wrong when i plug in the v(0)

OpenStudy (anonymous):

alright so when you plug in v(0) what did you get before adding like terms

OpenStudy (anonymous):

or before adding and subtracting... your you should get i+k=0-2j+0+c

OpenStudy (anonymous):

so then it is -2??

OpenStudy (anonymous):

meaning that c=i+2j+k

OpenStudy (anonymous):

so it should be +2

OpenStudy (anonymous):

k i got it right now

OpenStudy (anonymous):

but how did u know to do that or why did u do that??

OpenStudy (anonymous):

yeah when you put in v(0) you are basically trying to find c

OpenStudy (anonymous):

so for r(t) the i part do i have to do the same thing?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

because when you integrate what you have above you'll get another constant vector

OpenStudy (anonymous):

so will it bt +3

OpenStudy (anonymous):

be not bt..

OpenStudy (anonymous):

b/c r(0)= 1,1,1 right??

OpenStudy (anonymous):

i'm not sure i haven't integrated that far... these equations aare always sooo long lol hold on

OpenStudy (anonymous):

cos(2t)+t+1 i -sin(2t)+2t+1 j + 2/3(t^3)+t+1 k

OpenStudy (anonymous):

i got this and only got i wrong

OpenStudy (anonymous):

what did you get for your integration without using r(0)

OpenStudy (anonymous):

cos(2t)+t i - sin(2t)+2t j + 2/3(t^3) + t k

OpenStudy (anonymous):

you mean (-2t) for all your trig ops right/

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so if you do r(0)=i+j+k you should get i+j+k=i+0j+0k+c minus the i to the other side to get c by itself should get j+k =c

OpenStudy (anonymous):

what r(0) is basically saying is that when your t =0, your answer will be i+j+k

OpenStudy (anonymous):

y is j and k 0?

OpenStudy (anonymous):

so if you put 0 into the equation that you have above cos(-2t) you get 1

OpenStudy (anonymous):

(-sin(-2(0))+0)j = 0 sin(0) is 0

OpenStudy (anonymous):

so your c will be j+k if you put that back into the equation you'll get \[[\cos(-2t)+t]i+[-\sin(-2t)+2t+1]j+[\frac{2}{3}t^3+t+1]\]

OpenStudy (anonymous):

so for i i should put cos(-2t)+t?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

yep the above should be the equation of the position vector

OpenStudy (anonymous):

k i got it right thanks a lot for your help i appreciate it

OpenStudy (anonymous):

when solving this problem you go through these steps 1) Integrate to get to the velocity or position vector function (DO NOT FORGET that integrating an indefinite integral produces a +C or a constant vector) 2)Use the given info r(0) of v(0) to find the constant(r(0)=i+j+k, it just means that when t=0, your equation equals i+j+k) 3)once finding c by doing the above put it into your r(t) or v(t)

OpenStudy (anonymous):

k i will

OpenStudy (anonymous):

yep no problem